OpenStudy (anonymous):

What are the solutions to the equation |2x 3| + 4 = 17?

OpenStudy (anonymous):

what is between the 2x and the 3

OpenStudy (anonymous):

a minus or plus

OpenStudy (anonymous):

The absolute value = what's within the brackets of the equation. But I'm sort of confused?

OpenStudy (anonymous):

Minus

OpenStudy (igreen):

There will be two solutions to this problem.

OpenStudy (igreen):

2x - 3 + 4 = 17 Subtract 4 to both sides: 2x - 3 = 13 Add 3 to both sides: 2x = 16 Divide 2 to both sides: x = 8 This is your first solution.

OpenStudy (anonymous):

2x-3+4=17?

OpenStudy (anonymous):

First you need to get the absolute value by itself so l 2x + 3l = 13 then you add the value of 3 so 2x+3=13= 2x=10 x=5 But since it is a absolute value you must subtract the 3 as well so 2x-3=13 2x=16 x=8

OpenStudy (igreen):

2x + 3 + 4 = 17 Subtract 4 to both sides: 2x + 3 = 13 Subtract 3 to both sides: 2x = 10 Divide 2 to both sides: x = 5 This is your 2nd solution.

OpenStudy (igreen):

Your equations were: 2x + 3 + 4 = 17 and 2x - 3 + 4 = 17 You just change the sign inside of the absolute value, and solve both equations for x.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Taylor3333